solve the system of equations by graghing
x+2y=2
2x+4y=36

Answer by mananth(16946)

x+2y=2...........1
2x+4y=36...........2

equation1
when x= 0 y= 1 ( 0 , 1 )
when y = 0 x= 2 ( 2 , 0 )
Equation 2
when x= 0 y= 9 ( 0 , 9 )
when y = 0 x= 18 ( 18 , 0 )

You will observe that the graphs have same slope and different intercepts.
So the lines are parallel to each other.
m.ananth@hotmail.ca